# Segre forms of singular metrics on vector bundles and Lelong numbers

### v1

@@ 4. Generalized cycles and quasi-cycles
Using (4.2), it follows as in [5, Lemma 3.2] that if $V\hookrightarrow Y$ is an analytic subspace, and $\mu\in{\mathcal{QZ}}(Y)$, then ${\bf 1}_{V}\mu\in{\mathcal{QZ}}(Y)$. Indeed, if
@@
where $\tau_{j}:W_{j}\to Y$, then
(4.6) ${\bf 1}_{V}\mu=\sum_{\gamma_{j}(W_{j})\subset V}(\tau_{j})_{*}\gamma_{j}.$
@@
The proof is precisely the same as for the analogous statement for ${\mathcal{GZ}}(Y)$, [5, Proposition 3.6].
@@
If $\mu\in{\mathcal{QZ}}(Y)$ has support at $x\in Y$, then since it has order 00 it must be a number times $[x]$. In view of Proposition 4.4 this number is a finite sum of numbers $\alpha=\int\tau_{*}\gamma$, where $\tau(W)=\{x\}$. Now
$\alpha=\int_{Y}\tau_{*}\gamma=\int_{W}\gamma$
@@
4.1. Monge-Ampère type products as homomorphisms on ${\mathcal{QZ}}(Y)$
Let $u$ be a qas function on the complex manifold $Y$. Let $\pi\colon Y^{\prime}\to Y$ be a modification which principalizes the ideals associated to $\pi^{*}u$, i.e., such that in any local representation (2.5), $f$ defines a principal ideal. Such a modification indeed exists. In the case of integral analytic singularities, note first that we may always choose $c=1$ in (2.5). If we then have two different such representations with $f$ and $f^{\prime}$, then $\log(|f^{\prime}|^{2}/|f|^{2})$ is locally bounded, i.e., $|f^{\prime}|\sim|f|$. If we let ${\mathcal{J}}$ denote the integral closure of the ideal generated by $f$, we thus obtain a well-defined coherent ideal sheaf on $Y$. By blowing up ${\mathcal{J}}$ (followed by a resolution of singularities if one wants $Y^{\prime}$ to be smooth), one obtains such a modification. For the case when the analytic singularities are rational or real, see Remark 4.11.
We may thus assume that there is a section $\sigma$ of a holomorphic line bundle ${\mathcal{L}}\to Y^{\prime}$ such that if we equip ${\mathcal{L}}$ with some Hermitian norm, then
@@ 4.1. Monge-Ampère type products as homomorphisms on ${\mathcal{QZ}}(Y)$
where $\nu=-c_{1}({\mathcal{L}})+a$ is a global current on $Y^{\prime}$ of the type that occurs in (4.1).
@@
Note that $\nu$ does not depend on the choice of Hermitian norm on ${\mathcal{L}}$. One can in fact incorporate $a$ into the norm on ${\mathcal{L}}$, i.e., choose a norm $\|\cdot\|^{2}$ on ${\mathcal{L}}$ (with empty unbounded locus) so that $\pi^{*}u=\log\|\sigma\|^{2}$. ∎
@@
Assume that $\gamma$ is a product on $W$ as in (4.1) and $f$ is a non-trivial holomorphic function. Then $(\log|f|^{2})\gamma$ has locally finite mass, and $dd^{c}((\log|f|^{2})\gamma)=[\operatorname{div}f]\wedge\gamma.$ If $u_{j}$ are smooth and decrease to $\log|f|^{2}$, then $u_{j}\gamma\to(\log|f|)\gamma$.
@@
The first statement follows from [13, Theorem III.4.5] since locally in $Y^{\prime}$ all $b_{j}$ in $\gamma$ are $b_{j}=b_{j}^{\prime}+a_{j}$ where $b_{j}^{\prime}$ are psh and bounded and $a_{j}$ are smooth. The second and third one follow from [13, Proposition III.4.9]. ∎
@@
Let $u$ be a qas function on $Y$ with singular set $Z$. If $\mu\in{\mathcal{QZ}}(Y)$, then $u{\bf 1}_{Y\setminus Z}\mu$ is a measure with locally finite mass and
$[dd^{c}u]\wedge\mu:=dd^{c}(u{\bf 1}_{Y\setminus Z}\mu)$
@@
By linearity we can assume that $\mu=\tau_{*}\gamma$,  $\tau\colon W\to Y$. If the image of $\tau$ is contained in $Z$, then ${\bf 1}_{Y\setminus Z}\mu=0$. Otherwise $\tau^{*}u\not\equiv 0$ is qas on $W$ by Lemma 2.1. Then
@@
since $\tau^{-1}Z$ has positive codimension on $W$, cf. (4.2). In view of the discussion above we may assume, after possibly performing a modification of $W$, that $\tau^{*}u=\log|\sigma|^{2}+a$ on $W$, cf. (4.7), where $a$ is locally bounded.
Since $u$ is upper semi-continuous there is a sequence $u_{j}$ of smooth functions that decreases to $u$. By Lemma 4.7, $\tau^{*}u_{j}\gamma$ tends to the locally finite measure $\tau^{*}u\gamma$. It follows that
@@
is a locally finite measure. Moreover, $dd^{c}(u\mu)=\tau_{*}(dd^{c}(\tau^{*}u\gamma))$ so it is enough to see that $dd^{c}(\tau^{*}u\gamma)$ is in ${\mathcal{QZ}}(W)$.
By (4.8) on $W$, and Lemma 4.7,
@@
The last term in (4.9) is of the form (4.1). Notice that $[\operatorname{div}\sigma]$ is a sum of terms $m_{\ell}[D_{\ell}]$, where $m_{\ell}$ are integers and $D_{\ell}$ are divisors in $W$. In view of Lemma 4.3 (ii), the first term on the right hand side of (4.9) is therefore a finite sum of terms $m_{\ell}(i_{\ell})_{*}(i_{\ell}^{*}\gamma)$, where $i_{\ell}\colon D_{\ell}\to W$, and each of them is in ${\mathcal{QZ}}(W)$. ∎
@@
(4.10) $[dd^{c}u]^{0}\wedge\mu=\mu,\quad[dd^{c}u]^{k+1}\wedge\mu=[dd^{c}u]\wedge[dd^{c}u]^{k}\wedge\mu,\quad k=0,1,2,....$
If we let [^3] If $u=\log|\sigma|^{2}$, then $M^{u}$ is denoted by $M^{\sigma}$ in, e.g., [5]. We hope this will not cause any confusion.
(4.11) $\langle dd^{c}u\rangle^{k}\wedge\mu:={\bf 1}_{X\setminus Z}[dd^{c}u]^{k}\wedge\mu,\quad M^{u}_{k}\wedge\mu:={\bf 1}_{Z}[dd^{c}u]^{k}\wedge\mu,$
we see that, e.g.,
(4.12) $[dd^{c}u]^{k+1}\wedge\mu=dd^{c}(u\langle dd^{c}u\rangle^{k}\wedge\mu).$
@@
(4.13) $[dd^{c}\tau^{*}u]^{k}\wedge\gamma=M_{k}^{\tau^{*}u}\wedge\gamma+\langle dd^{c}\tau^{*}u\rangle^{k}\wedge\gamma=[\operatorname{div}\sigma]\wedge\nu^{k-1}\wedge\gamma+\nu^{k}\wedge\gamma.$
@@
If $\tau\colon Y^{\prime}\to Y$ is a proper mapping $\mu\in{\mathcal{QZ}}(Y^{\prime})$, then
(4.14) $[dd^{c}u]^{k}\wedge\tau_{*}\mu=\tau_{*}\big{(}[dd^{c}\tau^{*}u]^{k}\wedge\mu\big{)},\quad\langle dd^{c}u\rangle^{k}\wedge\tau_{*}\mu=\tau_{*}\big{(}\langle dd^{c}\tau^{*}u\rangle^{k}\wedge\mu\big{)},\\ \quad M_{k}^{u}\wedge\tau_{*}\mu=\tau_{*}\big{(}M_{k}^{\tau^{*}u}\wedge\mu\big{)},\quad k=0,1,2,\ldots.$
@@
If $u$ is $\mathbb{Q}$-qas or $\mathbb{R}$-qas, we claim that one may indeed also in this case find a modification $\pi:Y^{\prime}\to Y$ such that in the local representations (2.5) of $\pi^{*}u$, the ideal generated by $f$ is principal. It then follows readily that the results in this section hold for such $u$ when ${\mathcal{QZ}}$ is replaced by ${\mathcal{QZ}}_{\mathbb{Q}}$ or ${\mathcal{QZ}}_{\mathbb{R}}$.
To prove the claim, we first claim that there is a well-defined minimal modification $\pi^{\prime\prime}:Y^{\prime\prime}\to Y$, where $Y^{\prime\prime}$ is normal and such that the ideal generated by $f$ in the local representation (2.5) of $\pi^{*}u$ is principal. Given such a modification, one obtains $\pi:Y^{\prime}\to Y$ by possible taking the corresponding reduced space, and taking a resolution of singularities $\pi^{\prime}:Y^{\prime}\to Y^{\prime\prime}$.
If $u$ is ${\mathbb{Q}}$-qas, then one readily verifies that $Y^{\prime\prime}$ may be defined locally by blowing up the ideal generated by $f$ in the local representation (2.5) of $u$ followed by the normalization. Given two representations (2.5) with $c,f$ and $c^{\prime},f^{\prime}$, they will be such that $|f^{\prime}|\sim|f|^{c/c^{\prime}}$, where $c/c^{\prime}\in{\mathbb{Q}}$, and then it follows by the following well-known facts that the above procedure is well-defined: If ${\mathcal{J}}\subset{\mathcal{O}}_{Y}$ is a coherent ideal, and $\bar{{\mathcal{J}}}$ denotes its integral closure, and $\operatorname{Bl}^{+}_{\mathcal{J}}Y$ denotes the normalization of the blowup of ${\mathcal{J}}$, then $\operatorname{Bl}^{+}_{{\mathcal{J}}^{k}}Y=\operatorname{Bl}^{+}_{\mathcal{J}}Y=\operatorname{Bl}^{+}_{\bar{{\mathcal{J}}}}Y$ for $k\geq 1$.
If $u$ is ${\mathbb{R}}$-qas, then given two representations (2.5) with $c,f$ and $c^{\prime},f^{\prime}$, then if $x\in Z(f)=Z(f^{\prime})$, the Lelong number $\nu(x,u)$ is a positive integer times $c$ and an positive integer times $c^{\prime}$, and hence, $c/c^{\prime}\in{\mathbb{Q}}$, and one may reduce to the rational case. ∎
4.2. Action of smooth Segre and Chern forms
@@
defined by suitable compositions, cf. (2.4). Since these operators are just multiplication by $s_{k}(E,h)$ and $c_{k}(E,h)$, in particular we have that ${\bf c}_{0}(E,h)={\bf s}_{0}(E,h)=I$, the identity on ${\mathcal{QZ}}(X)$.
@@ 4.3. The submodule ${\mathcal{QZ}}_{0}(Y)$
Let $L$ be a holomorphic line bundle over $W$ with two different metrics $e^{b}$ and $e^{b^{\prime}}$, where $b^{\prime}$ is assumed to be smooth, and $b$ is qpsh and locally bounded. Then $v=b-b^{\prime}$ is a locally bounded qpsh function, and thus $e^{v}$ may be considered as a metric on the trivial line bundle. Let $\beta:=dd^{c}b^{\prime}-dd^{c}b=dd^{c}v$. Again we have a mapping ${\mathcal{QZ}}(W)\to{\mathcal{QZ}}(W)$,  $\mu\mapsto\beta\wedge\mu$. We say that any of these two forms $\beta$ is a beta form.
@@
We let ${\mathcal{QZ}}_{0}(Y)$ be the submodule of ${\mathcal{QZ}}(Y)$ generated by all currents of the form $\tau_{*}\mu^{\prime}$, $\tau\colon W\to Y$, where $\mu^{\prime}$ is in the image of one of the mappings $\beta$ above.
Thus ${\mathcal{QZ}}_{0}(Y)$ is generated by all $\tau_{*}(\beta\wedge\gamma)$ where $\beta$ is a beta form and $\gamma$ is a product of the form (4.1). For each such form there is a locally bounded function $v$ such that
(4.19) $dd^{c}\tau_{*}(v\wedge\gamma)=\tau_{*}(\beta\gamma).$
@@
If $\tau\colon Y^{\prime}\to Y$ is a proper mapping, then $\tau_{*}\colon{\mathcal{QZ}}_{0}(Y^{\prime})\to{\mathcal{QZ}}_{0}(Y)$. In view of (4.6), for any analytic subspace $V\subset Y$ we have that
@@
Remark 4.15.
Assume that $\mu\in{\mathcal{QZ}}_{0}(Y)$, and $\mu$ support on the analytic subspace $V\subset Y$. It then follows from Proposition 4.4 and (4.19) that there is a current $w$ of order 00 which is the pushforward of a current on $V$ such that $dd^{c}w=\mu$.
In particular, if $\mu\in{\mathcal{QZ}}_{0}(Y)$ has support at a point $x$ therefore
[…]
and since $\mu=\alpha[x]$, cf. Example 4.5, $\mu=0$. ∎
@@
$[dd^{c}u]^{k}\colon{\mathcal{QZ}}_{0}(Y)\to{\mathcal{QZ}}_{0}(Y),\quad k=0,1,2,\ldots.$
@@
$[dd^{c}u]^{k}-[dd^{c}u^{\prime}]^{k}\colon{\mathcal{QZ}}(Y)\to{\mathcal{QZ}}_{0}(Y),\quad k=0,1,2,\ldots.$
@@
Recall that if $\mu$ has support in $Z$, then $[dd^{c}u]\wedge\mu=0$. Otherwise, assume that $\mu=\tau_{*}(\beta\wedge\gamma)$ where $\gamma$ is product as in (4.1) and $\beta$ is a beta form. As in the proof of Proposition 4.8 we can therefore assume, cf. (4.9), that
(4.22) $[dd^{c}u]\wedge\mu=\tau_{*}\big{(}[\operatorname{div}\sigma]\wedge\beta\wedge\gamma+\nu\wedge\beta\wedge\gamma\big{)}.$
If $i\colon D\to W$, then $i^{*}(\beta\wedge\gamma)=i^{*}\beta\wedge i^{*}\gamma$ and thus $[\operatorname{div}\sigma]\wedge\beta\wedge\gamma$ is in ${\mathcal{QZ}}_{0}(W)$. Thus (i) holds for $k=1$ and hence for a general $k$ by induction.
By (i) and a simple induction over $k$ it is enough to prove (ii) for $k=1$. Let us assume that $\mu=\tau_{*}\gamma$, $\tau\colon W\to Y$. After a suitable further modification we can assume that $\tau^{*}u=\log|\sigma|^{2}+a$ and $\tau^{*}u^{\prime}=\log|\sigma^{\prime}|^{2}+a^{\prime}$ in $W$ where $a$ and $a^{\prime}$ are locally bounded. The assumption in (ii) implies that $[\operatorname{div}\sigma^{\prime}]=[\operatorname{div}\sigma]$. If $a^{\prime\prime}$ is some smooth metric on the line bundle corresponding to those divisors, then $dd^{c}(a-a^{\prime})=dd^{c}(a-a^{\prime\prime})-dd^{c}(a^{\prime}-a^{\prime\prime})=\beta-\beta^{\prime}$ where both $\beta$ and $\beta^{\prime}$ are beta forms on $W$. Therefore,

### v2

@@ 4. Generalized cycles and quasi-cycles
Using (4.2), it follows, cf. [5, Lemma 3.2], that if $V\hookrightarrow Y$ is an analytic subspace and $\mu\in{\mathcal{QZ}}(Y)$, then ${\bf 1}_{V}\mu\in{\mathcal{QZ}}(Y)$. Indeed, if
@@
where $\tau_{j}\colon W_{j}\to Y$, then
(4.6) ${\bf 1}_{V}\mu=\sum_{\tau_{j}(W_{j})\subset V}(\tau_{j})_{*}\gamma_{j}.$
@@
The proof is the same as for the analogous statement for ${\mathcal{GZ}}(Y)$, [5, Proposition 3.6].
@@
If $\mu\in{\mathcal{QZ}}(Y)$ has support at $x\in Y$, then since it has order 0 it must be a number times $[x]$, the point evaluation current at $x$. In view of Proposition 4.4 this number is a finite sum of numbers $\alpha=\int\tau_{*}\gamma$, where $\tau(W)=\{x\}$. Now
$\alpha=\int_{Y}\tau_{*}\gamma=\int_{W}\gamma,$
@@
4.1. Monge-Ampère type products as endomorphisms on ${\mathcal{QZ}}(Y)$
Assume that $u$ is a qas function on $Y$. Let $f$ be the tuple in a local representation (2.5) in $U\subset Y$ (with $c=1$) and let $\pi\colon U^{\prime}\to U$ be the normalization of the blowup of $Y$ along (the ideal generated by) $f$. If $\sigma$ is a section of the line bundle ${\mathcal{L}}\to U^{\prime}$ defining the exceptional divisor in $U^{\prime}$, then $\pi^{*}f=\sigma f^{\prime}$ where $f^{\prime}$ is a non-vanishing tuple of sections of ${\mathcal{L}}^{-1}$. If $\tilde{f}$ is the tuple from another representation (2.5) in $U$, then $|\tilde{f}|\sim|f|$ and therefore this modification is also the normalization of the blowup along $\tilde{f}$. This follows from the minimality property of such a modification. Therefore we get a global normal modification $p\colon Y^{\prime}\to Y$  such that any $f$ occurring in some local (2.5) is principal. (If wanted, by taking a further modification, one can assume that $Y^{\prime}$ is smooth but this is not necessary.) We thus get a global section $\sigma$ of a holomorphic line bundle ${\mathcal{L}}\to Y^{\prime}$ such that, if we equip $\mathcal{L}$} with some Hermitian norm,
@@ 4.1. Monge-Ampère type products as endomorphisms on ${\mathcal{QZ}}(Y)$
where $\nu=-c_{1}({\mathcal{L}})+a$ is a global current on $Y^{\prime}$ of the form (4.1).
@@
Note that $\nu$ does not depend on the choice of Hermitian norm on $\mathcal{L}$}. One can in fact incorporate $a$ into the norm on $\mathcal{L}$}, i.e., choose a qas norm $\|\cdot\|^{2}$ on $\mathcal{L}$} (with empty degeneracy locus) so that $\pi^{*}u=\log\|\sigma\|^{2}$. ∎
@@
Assume that $\gamma$ is a product on $W$ as in (4.1) and $f$ is a non-trivial holomorphic function. Then $(\log|f|^{2})\gamma$ has locally finite mass, and $dd^{c}((\log|f|^{2})\gamma)=[\operatorname{div}f]\wedge\gamma.$ If $u_{j}$ are smooth and decrease to $\log|f|^{2}$, then $u_{j}\gamma\to(\log|f|^{2})\gamma$.
@@
The first statement follows from [13, Theorem III.4.5] since locally in $Y^{\prime}$ all $b_{j}$ in $\gamma$ are $b_{j}=b_{j}^{\prime}+a_{j}$ where $b_{j}^{\prime}$ are psh and bounded and $a_{j}$ are smooth. The second statement follows from [13, Proposition III.4.9] and the third one from dominated convergence. ∎
@@
Let $u$ be a qas function on $Y$ with degeneracy locus $Z$. If $\mu\in{\mathcal{QZ}}(Y)$, then $u{\bf 1}_{Y\setminus Z}\mu$ has order zero and locally finite mass and
(4.9) $dd^{c}u]\wedge\mu:=dd^{c}(u{\bf 1}_{Y\setminus Z}\mu)$
@@
We interpret the right hand side of (4.9) as 0 if $\mu$ has support in $Z$. In particular, $dd^{c}u]\wedge\mu=0$ for all $\mu$ if $u\equiv$-$\infty$.
@@
By linearity we can assume that $\mu=\tau_{*}\gamma$, $\tau\colon W\to Y$. If the image of $\tau$ is contained in $Z$, then ${\bf 1}_{Y\setminus Z}\mu=0$. Otherwise $\tau^{*}u\not\equiv$-$\infty$ is qas on $W$ by Lemma 2.1. Then
@@
since $\tau^{-1}Z$ has positive codimension on $W$, cf. (4.2). In view of the discussion above we may assume, after possibly taking a modification of $W$, that $\tau^{*}u=\log|\sigma|^{2}+a$ on $W$, cf. (4.7), where $a$ is locally bounded.
Since $u$ is upper semi-continuous there is a sequence $u_{j}$ of smooth functions that decreases to $u$. By Lemma 4.7, $\tau^{*}u_{j}\gamma$ tends to the $\tau^{*}u\gamma$ that has locally finite mass. It follows that
@@
is a locally finite measure. Moreover, $dd^{c}(u\mu)=\tau_{*}(dd^{c}(\tau^{*}u\gamma))$ so it is enough to see that $dd^{c}(\tau^{*}u\gamma)$ is in ${\mathcal{QZ}}(W)$. By (4.8) on $W$, and Lemma 4.7,
@@
The last term in (4.10) is of the form (4.1). Notice that $\operatorname{div}\sigma]$ is a sum of terms $m_{\ell}[D_{\ell}]$, where $m_{\ell}$ are integers and $D_{\ell}$ are divisors in $W$. In view of Lemma 4.3 (ii), the first term on the right hand side of (4.10) is therefore a finite sum of terms $m_{\ell}(i_{\ell})_{*}(i_{\ell}^{*}\gamma)$, where $i_{\ell}\colon D_{\ell}\to W$, and each of them is in ${\mathcal{QZ}}(W)$. ∎
@@
(4.11) $dd^{c}u]^{0}\wedge\mu=\mu,\quad[dd^{c}u]^{k+1}\wedge\mu=[dd^{c}u]\wedge[dd^{c}u]^{k}\wedge\mu,\quad k=0,1,2,.$...
Let [^3] If $u=\log|\sigma|^{2}$, then $M^{u}$ is denoted by $M^{\sigma}$ in, e.g., [5]. We hope this will not cause any confusion.
(4.12) $\langle dd^{c}u\rangle^{k}\wedge\mu:={\bf 1}_{X\setminus Z}[dd^{c}u]^{k}\wedge\mu,\quad M^{u}_{k}\wedge\mu:={\bf 1}_{Z}[dd^{c}u]^{k}\wedge\mu.$
We have, cf. (4.9),
(4.13) $dd^{c}u]^{k+1}\wedge\mu=dd^{c}(u\langle dd^{c}u\rangle^{k}\wedge\mu).$
@@
(4.14) $dd^{c}\tau^{*}u]^{k}\wedge\gamma=M_{k}^{\tau^{*}u}\wedge\gamma+\langle dd^{c}\tau^{*}u\rangle^{k}\wedge\gamma=[\operatorname{div}\sigma]\wedge\nu^{k-1}\wedge\gamma+\nu^{k}\wedge\gamma.$
@@
If $\tau\colon Y^{\prime}\to Y$ is a proper mapping and $\mu\in{\mathcal{QZ}}(Y^{\prime})$, then
(4.15) $dd^{c}u]^{k}\wedge\tau_{*}\mu=\tau_{*}\big([dd^{c}\tau^{*}u]^{k}\wedge\mu\big),\quad\langle dd^{c}u\rangle^{k}\wedge\tau_{*}\mu=\tau_{*}\big(\langle dd^{c}\tau^{*}u\rangle^{k}\wedge\mu\big),\\ \quad M_{k}^{u}\wedge\tau_{*}\mu=\tau_{*}\big(M_{k}^{\tau^{*}u}\wedge\mu\big),\quad k=0,1,2,\ldots.$
@@
We claim that if $u$ is $\mathbb{Q}$-$\mathbb{R}-qas$  then there is a modification $\pi\colon Y^{\prime}\to Y$ and a section $\sigma$ of a line bundle ${\mathcal{L}}\to Y^{\prime}$ such that locally on $Y^{\prime}$, $\pi^{*}u=c\log|\sigma|^{2}+a$, where $a$ is qpsh and $c>0$ is rational/real. Notice that $c$ cannot be globally constant in general. For instance, let $u$ be a potential to $\sum_{k}(1/k)[z=k]$ in the complex plane.
To see the claim, first recall that the normalization $p\colon U^{\prime}\to U$ of the blowup of $U$ along the ideal $\mathcal{J}$} generated by $f$ is the same as the blow up of the ideal ${\mathcal{J}}^{k}$ for any $k\geq 1$. If $f$ and $\hat{f}$ occur in two local representations (2.5) then $c^{\prime}\log|f^{\prime}|=c\log|f|+{\mathcal{O}}(1)$. Therefore $|f^{\prime}|^{k^{\prime}}\sim|f|^{k}$ for some integers $k,k^{\prime}\geq 1$. This is obvious if $u$ is $\mathbb{Q}$-qas so that $c,c^{\prime}\in\mathbb{Q}$. In the real case this is clear if $f$ and $f^{\prime}$ are functions, and one is reduced to this case in a suitable modification where $f$ and $f^{\prime}$ are principal. Thus $Y^{\prime}$ is well-defined locally and hence globally. Now the claim follows if $\sigma$ is a global section that defines the exceptional divisor. ∎
4.2. Action of Segre and Chern forms associated with a smooth metric
@@
defined by suitable compositions, cf. (2.4). Since these operators are just multiplication by $s_{k}(E,h)$ and $c_{k}(E,h)$, in particular ${\bf c}_{0}(E,h)={\bf s}_{0}(E,h)=I$, the identity on ${\mathcal{QZ}}(X)$.
@@ 4.3. The submodule ${\mathcal{QZ}}_{0}(Y)$
Let $L$ be a holomorphic line bundle over $W$ with two different metrics $e^{b}$ and $e^{b^{\prime}}$, where $b^{\prime}$ is assumed to be smooth, and $b$ is qpsh and locally bounded. Then $v=b-b^{\prime}$ is a locally bounded qpsh function. Let $\beta:=dd^{c}b^{\prime}-dd^{c}b=dd^{c}v$. Again we have a mapping ${\mathcal{QZ}}(W)\to{\mathcal{QZ}}(W)$, $\mu\mapsto\beta\wedge\mu$. We say that any of these two forms $\beta$ is a beta form.
@@
We let ${\mathcal{QZ}}_{0}(Y)$ be the submodule of ${\mathcal{QZ}}(Y)$ generated by all currents of the form $\tau_{*}\mu^{\prime}$, $\tau\colon W\to Y$, where $\mu^{\prime}$ is in the image of one of the mappings above.
Thus ${\mathcal{QZ}}_{0}(Y)$ is generated by all $\tau_{*}(\beta\wedge\gamma)$ where $\beta$ is a beta form and $\gamma$ is a product as in (4.1). For each such form there is a locally bounded function $v$ such that
(4.20) $dd^{c}\tau_{*}(v\gamma)=\tau_{*}(\beta\wedge\gamma).$
@@
If $\tau\colon Y^{\prime}\to Y$ is a proper mapping, then $\tau_{*}\colon{\mathcal{QZ}}_{0}(Y^{\prime})\to{\mathcal{QZ}}_{0}(Y)$. In view of (4.6), for any analytic subspace $V\subset Y$,
@@
Proposition 4.15.
If $\mu\in{\mathcal{QZ}}_{0}(Y)$ has bidegree $(k,k)$ and support on the analytic subspace $V\subset Y$, then $\mu=0$ if $k-1<{\rm codim\,}V$.
In particular, if $\mu\in{\mathcal{QZ}}_{0}(Y)$ has support at a point, then $\mu=0$.
Proof.
It follows from Proposition 4.4 and (4.20) that there is a current $w=i_{*}w^{\prime}$, where $i\colon V\to X$, such that $dd^{c}w=\mu$. If $\mu$ has bidegree $(k,k)$, then $w$ has bidegree (k1,k1)(k-1,k-1). Thus $w^{\prime}$ has bidegree $(k-1-{\rm codim\,}V,k-1-{\rm codim\,}V)$ and hence it vanishes if $k-1<{\rm codim\,}V$. ∎
@@
$dd^{c}u]^{k}\colon{\mathcal{QZ}}_{0}(Y)\to{\mathcal{QZ}}_{0}(Y),\quad k=0,1,2,\ldots.$
@@
$dd^{c}u]^{k}-[dd^{c}u^{\prime}]^{k}\colon{\mathcal{QZ}}(Y)\to{\mathcal{QZ}}_{0}(Y),\quad k=0,1,2,\ldots.$
@@
Recall that if $\mu$ has support in $Z$, then $dd^{c}u]\wedge\mu=0$. Otherwise, assume that $\mu=\tau_{*}(\beta\wedge\gamma)$ where $\gamma$ is product as in (4.1) and $\beta$ is a beta form. As in the proof of Proposition 4.8 we can therefore assume, cf. (4.10), that
(4.22) $dd^{c}u]\wedge\mu=\tau_{*}\big([\operatorname{div}\sigma]\wedge\beta\wedge\gamma+\nu\wedge\beta\wedge\gamma\big).$
If $i\colon D\to W$, then $i^{*}(\beta\wedge\gamma)=i^{*}\beta\wedge i^{*}\gamma$ and thus $\operatorname{div}\sigma]\wedge\beta\wedge\gamma$ is in ${\mathcal{QZ}}_{0}(W)$. Thus (i) holds for $k=1$ and hence for a general $k$ by induction.
By (i) and a simple induction over $k$ it is enough to prove (ii) for $k=1$. Let us assume that $\mu=\tau_{*}\gamma$, $\tau\colon W\to Y$. After a suitable further modification we can assume that $\tau^{*}u=\log|\sigma|^{2}+a$ and $\tau^{*}u^{\prime}=\log|\sigma^{\prime}|^{2}+a^{\prime}$ in $W$ where $a$ and $a^{\prime}$ are locally bounded. The assumption in (ii) implies that $\operatorname{div}\sigma^{\prime}]=[\operatorname{div}\sigma]$. If $a^{\prime\prime}$ is some smooth metric on the dual of the line bundle corresponding to this divisor, then $dd^{c}(a-a^{\prime})=dd^{c}(a-a^{\prime\prime})-dd^{c}(a^{\prime}-a^{\prime\prime})=\beta-\beta^{\prime}$ where both $\beta$ and $\beta^{\prime}$ are beta forms on $W$. Therefore,